Multiple choice
1. Suppose product price is fixed at BAM 24; MR = MC at Q = 200; AFC = BAM 6; AVC = BAM 16.
What do you advise this firm to do?
a. Increase output.
b. Shut down operations.
c. Stay at the current output; the firm is earning a profit of BAM 400.
d. Stay at the current output; the firm is losing BAM 200.

2. If a potato farmer expands output, he finds that the increase in total revenue is less than the increase in total costs. This means that:
a. profit is being maximized.
b. he should not have expanded output.
c. he should produce even more output.
d. the firm is wasting resources.
e. the farmer should go out of business.
1
Expert's answer
2015-06-29T00:00:44-0400
Multiple choice
1. Suppose product price is fixed at BAM 24; MR = MC at Q = 200; AFC = BAM 6; AVC = BAM 16.
As MR = MC and TP = (P - ATC)*Q = (24 - (16+6))*200 = 400, the I advise this firm to:
c. Stay at the current output; the firm is earning a profit of BAM 400.
2. If a potato farmer expands output, he finds that the increase in total revenue is less than the increase in total costs. This means that:
b. he should not have expanded output.
Because his profit decreased